# Trends and Reversion in Financial Markets on Time Scales from Minutes to Decades

### v1

@@ 5 Synthesis and Discussion
To make $b_{T}$ and $c_{T}$ for different time scales $T$ comparable, the time horizon $N$ of the future return $R_{N}$ should scale with the time horizon $T$ of the trend used to predict it. We therefore rescale the coefficients $b_{T},c_{T}$ to $\tilde{b}_{T},\tilde{c}_{T}$ as discussed subsection 3.3:
$\tilde{b}_{T}=b_{T}\cdot\sqrt{T/60}\ ,\ \ \ \tilde{c}_{T}=c_{T}\cdot\sqrt{T/60}.$
@@
Figure 10: Regression coefficients from minutes to decades. Left: for time horizons of $T\leq 8$ days. Right: For time horizons $T\geq 8$ days. Although the minute-, daily- and monthly data sets cover different time periods, the results connect with each other within error margins.
@@
  • Time horizons from a few hours to several days: small trends tend to persistent ($\tilde{b}>0$), while strong trends tend to revert ($\tilde{c}<0$). In this range, $\tilde{b}$ and $\tilde{c}$ are small. The jump from minute- to daily data is still within estimation errors. The data are consistent with assuming that $c$ is constant and financial markets are approximately scale invariant. $\tilde{b}$ and $\tilde{c}$ seem to be universal, i.e., independent of the asset or time zone (within errors). This has been explained in [13] by modeling trends as purely social phenomena that arise in the social network of investors, which is assumed to be near a critical point.
@@
  • Time horizons from a few years to decades: $\tilde{b}$ and $\tilde{c}$ become weaker and switch sign at horizons of $\approx 2$ years, beyond which weak trends tend to revert and strong trends seem to persist. This could reflect economic cycles and long-term developments.
These results provide a detailed set of empirical observations that measure the scaling behavior of financial markets. Theoretical model of financial markets, such as [23, 24, 25, 26, 27, 28], must replicate them, as well as other stylized facts [11, 12].
@@
$V={1\over 2}\tilde{b}\phi^{2}-{1\over 4}\tilde{c}\phi^{4}+{1\over 6}\tilde{d}\phi^{6}+...,\ \ \ \text{where}\ \tilde{b},\tilde{d}\ll 1,\ \tilde{c}=-g^{*}.$
Here, $g^{*}>0$ is value of the the $\phi^{4}$ coupling constant at its renormalization group fixed point, to which it flows at large scales. The derivative $V^{\prime}(\phi)$ plays the role of the driving force in (5).
$\tilde{d}$ is an irrelevant coupling constant: it flows to zero at large scales $T>T_{1}$ for some $T_{1}$, but grows at small scales $T<T_{1}$, below which non-universal microscopic details of the molecules (corresponding to market details such as tick sizes and liquidity) become visible. This matches our observations, if we identify $T_{1}\approx 15-30$ minutes. $\tilde{b}$ is a relevant coupling constant: it flows to zero at small scales $T<T_{2}$ for some $T_{2}$, but grows at large scales $T>T_{2}$. In the range $T_{1}<T<T_{2}$, the model is near the critical point. There, it should be approximately scale invariant and characterized by universal critical exponents, such as Hurst exponents. Consistency with our observations requires a $T_{2}$ of roughly a few days.
@@ Acknowledgements
We would like to thank W. Breymann, A. Nikeghbali, and … for discussions, and Donat Maier for helping us to clean the intra-day and long-term data. This research is supported by grant no. PT00P2_206333 from the Swiss National Science Foundation.

### v2

@@ 5 Synthesis and Discussion
To make $b_{T}$ and $c_{T}$ for different time scales $T$ comparable, the time horizon $N$ of the future return $R_{N}$ should scale with the time horizon $T$ of the trend used to predict it. We therefore rescale the coefficients $b_{T},c_{T}$ to $\tilde{b}_{T},\tilde{c}_{T}$ as in (11):
$\tilde{b}_{T}=b_{T}\cdot\sqrt{T/60}\ ,\ \ \ \tilde{c}_{T}=c_{T}\cdot\sqrt{T/60}.$ (13)
@@
Figure 10: Regression coefficients from minutes to decades based on ansatz (1), combining (i) intraday data since 2010, (ii) daily data since 1990, and (iii) monthly data since 1692. Left: time horizons of $T\leq 8$ days. Right: time horizons $T\geq 8$ days. Although the three data sets cover different time periods, the results connect with each other within error margins.
@@
  • Time horizons from a few hours to several days: small trends tend to persist ($\tilde{b}>0$), while strong trends tend to revert ($\tilde{c}<0$). In this range, $\tilde{b}$ and $\tilde{c}$ are small. The jump from minute- to daily data is still within estimation errors. The data are consistent with assuming that $c$ is constant and financial markets are approximately scale invariant. $\tilde{b}$ and $\tilde{c}$ seem to be universal, i.e., independent of the asset or time zone (within errors). This has been explained in [20] by modeling trends as purely social phenomena that arise in the social network of investors, which is assumed to be near a critical point.
@@
  • Time horizons from a few years to decades: $\tilde{b}$ and $\tilde{c}$ become weaker and switch signs at horizons of $\approx 2$ years, beyond which weak trends tend to revert and strong trends seem to persist. This could reflect economic cycles and long-term developments.
These results provide a detailed set of empirical observations that measure the scaling behavior of financial markets. Theoretical model of financial markets, such as [17, 27, 28, 29, 30, 31, 32], must replicate them, as well as other stylized facts [18, 19].
@@
$V={1\over 2}\tilde{b}\phi^{2}-{1\over 4}\tilde{c}\phi^{4}+{1\over 6}\tilde{d}\phi^{6}+...,\ \ \ \text{where}\ \tilde{b},\tilde{d}\ll 1,\ \tilde{c}=-g^{*}.$ (14)
Here, it is assumed that ”price-minus-value” $\phi$ can be measured as today’s price minus a long-term average price, which is essentially the (rescaled) trend strength $\phi_{k}$ in (2) in the limit of large $T=2^{k}$, close to the correlation time. $g^{*}>0$ is the value of the $\phi^{4}$ coupling constant at its renormalization group fixed point, to which it tends (or, in renormalization group language, ”flows”) at large scales. The derivative $V^{\prime}(\phi)$ plays the role of the cubic driving force in (1).
$\tilde{d}$ is an irrelevant coupling constant in the renormalization group sense. This means that it shrinks towards zero at large scales $T>T_{1}$ for some $T_{1}$, but grows at small scales $T<T_{1}$, below which non-universal microscopic details of the molecules (corresponding to market details such as tick sizes and liquidity) become visible. This matches our observations, if we identify $T_{1}\approx 15-30$ minutes.
By contrast, $\tilde{b}$ is a relevant coupling constant in the renormalization group sense. This means that it shrinks to zero at small scales $T<T_{2}$ for some $T_{2}$, but grows at large scales $T>T_{2}$. In the range $T_{1}<T<T_{2}$, the model is near the critical point. There, it should be approximately scale invariant and characterized by universal critical exponents, such as Hurst exponents. Consistency with our observations requires a $T_{2}$ of roughly a few days.
@@ Acknowledgements
We would like to thank Henriette (formely Wolfgang) Breymann, Ashkan Nikeghbali, Thomas Lehérici, and Maximilan S. Janisch for discussions, Gisela Reichmuth for remarks on tick data, and Donat Maier for helping us to clean the intra-day and long-term data. This research is supported by grant no. PT00P2_206333 from the Swiss National Science Foundation.
Appendix: Even Regression Terms
For completeness, here we generalize the regression ansatz (1,9) to also include the quadratic and higher-order polynomial terms:
$R(t+1)=a+b\cdot\phi(t)+c_{2}\cdot\phi^{2}(t)+c\cdot\phi^{3}(t)+c_{4}\cdot\phi^{4}(t)+d\cdot\phi^{5}(t)+\epsilon(t+1).$ (15)
The regression results for daily data are shown in table 10. The first 2 colums are the results of table 3 with only linear and cubic terms. In the next 2 columns, we add the quadratic term. Its coefficient $c_{2}$ is not significant (t-statistics $<1$). It increases the in-sample $R^{2}$ (both for single and aggregated time scales), but it decreases the out-of-sample $R^{2}$. This is a classical sign of data overfitting.
The last 2 columns include all coefficients. The adjusted $R^{2}$ is similar as for the linear-plus-cubic regression. However, the $t$-statistics for the coefficients $b-d$ are generally smaller, which can be attributed to overfitting outlier events with $|t|\geq 2.5$. We conclude that the linear-plus-cubic model ist the most reliable one, although we cannot exclude that the response function may rise/fall more steeply than $\phi^{3}$ in the far tails.
Linear & cubic terms With quadratic term 5th order polynomial
Coeff. Value in % t-stat Value in % t-stat Value in % t-stat
$a$ $+1.33\pm 0.41$ $3.3$ $+1.45\pm 0.36$ $4.0$ $+1.59\pm 0.30$ $5.3$
$b$ $+1.29\pm 0.43$ $3.0$ $+1.30\pm 0.43$ $3.0$ $+0.33\pm 0.44$ $0.7$
$c_{2}$ $-0.15\pm 0.36$ $0.4$ $-0.50\pm 0.33$ $1.5$
$c$ $-0.62\pm 0.23$ $2.7$ $-0.63\pm 0.23$ $2.7$ $+0.36\pm 0.36$ $1.0$
$c_{4}$ $+0.07\pm 0.12$ $0.6$
$d$ $-0.15\pm 0.07$ $2.1$
$R^{2}$ in bp Single $T$ Aggr. Single $T$ Aggr. Single $T$ Aggr.
$R^{2}$ $1.31$ $4.91$ 1.34 5.12 1.58 6.31
$R^{2}_{adj}$ $1.03$ $3.98$ 0.84 3.50 0.96 4.04
Table 10: Extension of table 3 for daily returns to the general regression ansatz (15).
Table 11 adds the quadratic term to the regression for the 150-year set of monthly data without currencies. For trends horizons up to one year, the quadratic coefficient $c_{2}$ is clearly insignificant, as for daily data. For longer trends, it is negative with t-statistic 2.4, which is below our chosen significance threshold of $t=3$ for these long-term data. So we prefer not to add the quadratic regression term, although we cannot clearly rule it out either.
Trend horizons up to 1 year Trend horizons $2-16$ years
Coeff. Value (%) t-stat Value (%) t-stat Value (%) t-stat Value (%) t-stat
$a$ $+6.1\pm 0.9$ 7.3 $+6.7\pm 0.9$ 7.7 $+2.2\pm 0.7$ 3.0 $+3.3\pm 0.7$ 4.6
$b$ $+7.0\pm 1.4$ $5.0$ $+7.0\pm 1.4$ $5.0$ $-4.2\pm 0.9$ 4.8 $-3.9\pm 0.9$ 4.3
$c_{2}$ $-0.2\pm 0.7$ 0.3 $-1.1\pm 0.5$ 2.4
$c$ $-1.4\pm 0.4$ $3.3$ $-1.4\pm 0.4$ $3.3$ $+1.1\pm 0.2$ 5.2 $+0.9\pm 0.3$ 3.5
$R^{2}$ in bp Single $T$ Aggr. Single $T$ Aggr. Single Aggr. Single $T$ Aggr.
$R^{2}$ $11.0$ 15.4 $11.1$ 15.6 $5.0$ 8.5 7.5 12.0
$R^{2}_{adj}$ $5.7$ 8.4 $3.5$ 5.4 $2.2$ 4.4 3.3 6.1
Table 11: Extension of table 8 for monthly returns to the general regression ansatz (15).
Table 12 shows an analogous analysis for intraday returns, in which case we also add a term $e\cdot\text{sign}(\phi)$ to ansatz (15). The quadratic/quartic terms are not or only barely statistically significant, and they do not increase the adjusted R-squared, so we prefer to keep our regression model simple and drop these even terms as explanatory variables.
Trend horizons up to 16 minutes Trend horizons 1 hour to 1 day
Coeff. Value (%) t-stat Value (%) t-stat Value (%) t-stat Value (%) t-stat
$a$ $+0.017$ $0.8$ 0.055 2.8 $+0.003$ 0.1 $-0.008$ 0.3
$b$ $-0.912$ $13.6$0.890 13.3 $+0.132$ 2.9 $+0.159$ 3.5
$c_{2}$0.076 2.1 $-0.002$ 0.1
$c$ $+0.259$ $4.6$ 0.272 4.9 $-0.039$ 2.8 $-0.038$ 2.7
$c_{4}$ 0.006 0.7 $-0.001$ 0.2
$d$ $-0.038$ $4.2$0.042 4.6
$e$ $-0.282$ $12.5$0.287 12.5 $-0.071$ 3.9 $-0.067$ 3.7
$R^{2}$ in bp Single Aggr. Single Aggr. Single Aggr. Single Aggr.
$R^{2}$ $0.43$ 0.56 0.43 0.56 0.012 0.025 0.013 0.027
$R^{2}_{adj}$ $0.42$ 0.55 0.42 0.55 0.009 0.020 0.007 0.016
Table 12: Extension of tables 5,6 for intraday returns to the general regression ansatz (15).